OpenStudy (safia21):

A line parallel to y=-2x+3 that passes through (3,4)

OpenStudy (katrinakaif):

A line parallel would have the same slope as the equation given. The slope here is -2 . Plug it into your equation, and the following values and solve. (y-y1) = -2(x-x1) as (3,4) = (x1,y1)